### v1

# EaDeblur-GS: Event assisted 3D Deblur Reconstruction with Gaussian Splatting

Abstract
3D deblurring reconstruction techniques have recently seen significant advancements with the development of Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS). Although these techniques can recover relatively clear 3D reconstructions from blurry image inputs, they still face limitations in handling severe blurring and complex camera motion. To address these issues, we propose Event-assisted 3D Deblur Reconstruction with Gaussian Splatting (EaDeblur-GS), which integrates event camera data to enhance the robustness of 3DGS against motion blur. By employing an Adaptive Deviation Estimator (ADE) network to estimate Gaussian center deviations and using novel loss functions, EaDeblur-GS achieves sharp 3D reconstructions in real-time, demonstrating performance comparable to state-of-the-art methods.
Keywords:
3D Gaussian Splatting Event Cameras Neural Radiance Fields
1 Introduction
Reconstructing 3D scenes and objects from images has long been a research hotspot in computer vision and computer graphics. The advent of Neural Radiance Fields (NeRF) [4] has brought revolutionary advancements in photo-realistic novel view synthesis. Recently, The introduction of 3D Gaussian Splatting (3DGS) [1] has revolutionized 3D scene representation using Gaussian ellipsoids, achieving high-quality 2D rendering and faster speeds. However, in real-world applications, factors like camera shake and shutter speed often lead to image blurriness and inaccurate camera poses estimation, challenging the clear neural volumetric representation.
Therefore, several methods have been developed to handle blurriness in NeRFs and 3DGS. NeRF’s deblurring technology has undergone relatively early development, with Deblur-NeRF [3] being the first framework to tackle this issue, which utilized an analysis-by-synthesis approach to recover sharp NeRFs from blurry inputs. MP-NeRF[8] further enhances this by introducing a multi-branch fusion network and prior-based learnable weights to handle extremely blurry or unevenly blurred images. But the NeRF-based methods always consume extensive training time and rendering time. Some methods based on 3DGS has been developed because of its advantages in rendering and training speed. For instance, Wenbo Chen et al. [9] proposed Deblur-GS, which models the problem of motion blur as a joint optimization involving camera trajectory and time sampling. B. Lee etal. [2] assign the corrections on the rotation and scaling matrix of 3D gaussians by using a small MLP, enhancing the clarity of scenes reconstructed from blurred images. Nonetheless, these methods can only achieve clear 3D reconstruction results with mildly blurred input images. Consequently, additional data sources like event cameras have been introduced into 3D deblurring reconstruction. Event cameras, a bio-inspired sensor, offers high temporal resolution and has advantages in motion deblur. For NeRFs, EventNeRF [7] used event integration and color simulation for colored 3D representations, while Qi et al. [6] combined blurred images with event streams using innovative loss functions and the Event Double Integral (EDI) approach. For 3DGS, Yu proposed EvaGaussian[10], which enhanced 3DGS with assistance of event streams.
Despite achieving excellent performance in 3D deblurring reconstruction, the aforementioned methods still have limitations. For instance, RGB single-modality deblurring 3DGS and NeRF are often effective only in mildly blurred or simple camera motion scenarios. When input images are severely blurred, these methods usually fail to reconstruct 3D objects and can only produce relatively clear 2D renderings from certain angles. On the other hand, techniques using event cameras for NeRF 3D reconstruction are limited by NeRF’s training and rendering speed. Additionally, methods that incorporate event data into 3DGS deblurring reconstruction face challenges, such as inaccurate camera motion trajectory estimation.
To overcome these limitations, we propose a novel integration of event streams with 3DGS, namely Event-assisted 3D Deblur Reconstruction with Gaussian Splatting(EaDeblur-GS), aiming to guide the learning of better 3D Gaussian representations and tackle issues arising from blurry inputs.
Our contributions are summarized as follows:
  • We propose Event-assisted 3D Deblur Reconstruction with Gaussian Splatting (EaDeblur-GS), which incorporates blurry RGB images and event streams into Gaussian Splatting to recover sharp 3D representations.
  • We introduce a novel Adaptive Deviation Estimator(ADE) network to simulate the shaking motion during exposure accurately by estimating the deviations of Gaussians.
  • We comprehensively evaluate the proposed method and compare it with several baselines, demonstrating that our EaDeblur-GS achieves performance comparable to state-of-the-art methods while enabling real-time sharp image rendering.
2 Method
As shown in Fig.1, our method starts with the input of blurry RGB images and the corresponding event streams. We first employ the Event Double Integral (EDI) technique [5] to generate a set of latent sharp images. These images are then processed by COLMAP to achieve enhanced initial reconstruction and precise camera pose estimation. From this enhanced reconstruction, we create a set of 3D Gaussians. Subsequently, the positions of these Gaussians and the estimated camera poses are fed into our proposed ADE network to determine the positional deviations of the Gaussians. These adjusted 3D Gaussians are then projected onto each viewpoint, including corresponding latent viewpoints, to produce sharp image renderings. Furthermore, we integrate a Blurriness Loss to simulate the generation of authentic blurry images, and an Event Integration Loss to guide the Gaussian model in accurately capturing the true shapes of objects. This allows the model to learn precise 3D volume representations and achieve superior 3D reconstructions. The overview of our method is illustrated in Fig.1.
Figure 1: The overview of EaDeblur-GS
In the following sections, we detail how the ADE network estimates deviations, followed by an in-depth introduction to blurriness loss and event integration loss.
2.1 Adaptive Deviation Estimator
Motion blur caused by camera shake impedes sparse initial reconstruction due to unclear input images. To mitigate this issue, we employ the EDI method [5], which combines blurry images and corresponding event streams. The EDI model transforms a blurry image $I_{blur}$ into multiple sharp images ${I_{0},\cdots,I_{k}}$, under the assumption that the blurry image is the temporal average of latent sharp images, each represented by the accumulated events. Given a blurry image and its corresponding event bins $\{B_{k}\}_{k=1}^{b}$, the sharp image of viewpoint $I_{0}$ and each latent sharp image $I_{k}$ can be expressed as:
$I_{0}=\frac{(b+1)I_{blur}}{1+e^{\Theta\sum_{i=1}^{1}B_{i}}+\cdots+e^{\Theta\sum_{i=1}^{b}B_{i}}}.$ (1)
$I_{k}=\frac{(b+1)I_{blur}e^{\Theta\sum_{i=1}^{k}B_{i}}}{1+e^{\Theta\sum_{i=1}^{1}B_{i}}+\cdots+e^{\Theta\sum_{i=1}^{b}B_{i}}}.$ (2)
We then utilize COLMAP to estimate the poses of EDI-processed sharp images $\{\mathbf{P}_{k}\}_{k=0}^{b}=\mathrm{COLMAP}(\{I_{k}\}_{k=0}^{b})$, which also results in a more accurate initial sparse point cloud.
Inspired by [9] and [2], we represent latent poses as displacements of Gaussian ellipsoid centers $x_{j}$. To estimate the deviations of Gaussians in this context, we employ Adaptive Deviation Estimator network, which is consisted of a small Multi-Layer Perceptron (MLP) $\mathcal{F}\theta$ with three hidden linear layers and an embedding layer which transforms low-frequency positional information into high-frequency expressions. The ADE $\mathcal{F}\theta$ takes the EDI-predicted poses ${\mathbf{P}k}{k=0}^{b}$ and original positions of Gaussians $x_{j}$ to estimate deviations:
$\{(\delta x_{j}^{(i)}\}_{i=1}^{l}=\mathcal{F}_{\theta}\Big{(}\gamma(x_{j}),\gamma(P)\Big{)}$ (3)
where $l$ means the number of estimated latent poses, and $\delta x_{j}^{(i)}$ denotes the $i$-th predicted position offset of $j$-th Gaussian. We then obtain extra $l$ sets of 3D Gaussians $\{\{\hat{x}_{j}^{(i)}\}_{i=1}^{l}\}_{j=1}^{N_{G}}$ by adjusting the positions of the original 3D Gaussians, where $N_{G}$ is the number of Gaussians, $\hat{x}_{j}^{(i)}$ represents the position offset scaled by $\lambda_{p}$: $\delta x_{j}^{(i)}$ by $\lambda_{p}.\hat{x}_{j}^{(i)}=x_{j}+\lambda_{p}\delta x_{j}^{(i)}$. This method computes $l$ different sets of 3D Gaussians for each viewpoint, which can be rasterized to $l$ sharp latent images $\{I_{i}\}_{i=0}^{l}$. Additionally, we rasterize an image with original Gaussians positions as the image rendered from original viewpoint. During forward rendering process, the MLP network and multiple rendering are not necessary, ensuring real-time inference speed comparable to original 3D Gaussian.
2.2 Loss Functions
Blurriness loss. To model the motion blur process during exposure time, we take the average sum of rendered images as follows:
$\hat{I}_{blur}=\frac{1}{b+1}\sum_{i=0}^{b}I_{i},\quad I_{i}=\text{Rasterize}(\{G(\hat{x}_{j}^{(i)})\}_{j=1}^{N_{G}})$ (4)
where $I_{i}$ is a sharp image generated by the estimated corrections, $I_{blur}$ is the estimated blurry image. The blurriness loss is computed as the difference between estimated blurry image and input blurry image $I_{blur}$, combined with a D-SSIM loss as follows:
$L_{blur}=(1-\lambda_{D-SSIM})\|I_{blur}-\hat{I}_{blur}\|_{1}+\lambda_{D-SSIM}L_{D-SSIM}$ (5)
where we set $\lambda_{D-SSIM}=0.2$ for all experiments.
Event Integration Loss. Leveraging the high time-resolution event stream, we adopt an Event Integration Loss to guide the network in learning a fine-grained sharp 3D reconstructions. We integrate the events polarities between two input frames as follows:
[…]
where $\delta t$ is time interval between two input frames. The logarithm difference between the last rendered frame and first rendered frame of multiple rendered frames $\{I_{i}\}_{i=0}^{l}$ gives the estimated event integration maps $\tilde{\mathbf{E}}(t)$. The event integration loss is then:
$L_{event}=\|\mathbf{E}(t)-\widetilde{\mathbf{E}}(t)\|_{1}$ (7)
The final loss function is the combination of the blurriness loss and event integration loss as follows:
$L=L_{blur}+\lambda_{event}L_{event}$ (8)
3 Experiments
3.1 Comparisons and Results
To evaluate the effectiveness of using event data to aid in deblurring RGB images, we compare our method with the original Gaussian Splatting (GS), which utilizes blurry images as input and supervision. Since COLMAP fails with only blurry images, we estimate the initial point cloud using EDI-deblurred images.
We also compare our method with Deblurring 3D Gaussian Splatting(Deblurring-GS) to highlight the advantages of incorporating event data. Additionally, we benchmark our approach against Event Enhanced Neural Radiance Fields from Blurry Images(E2NeRF), the state-of-the-art method that leverages event data for deblurring RGB images.
Our evaluation metrics include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Frames Per Second (FPS). All methods are tested on the synthetic data from E2NeRF, which includes perfectly sharp images from the same viewpoint as blurry input images with corresponding event streams. The implementation details are provided in the supplementary materials.
The quantitative comparison results are presented in Table 1 and the qualitative comparison is also provided in Supplementary materials.
Table 1: Quantitative analysis. The results in tha table are the averages of six synthetic scenes from E2NeRF. We use bold to mark the best result.
Image Deblur GS Deblurring-GS E2NeRF Ours
PSNR 22.15 22.70 29.77 30.08
SSIM 0.878 0.8427 0.960 0.9366
FPS 160 160 2 160
Our results demonstrate that our method achieves performance comparable to the current state-of-the-art model in terms of PSNR and SSIM, while maintaining real-time rendering capabilities with noticeable FPS. The advantage of using event data for deblurring RGB images is further evidenced by comparing Deblurring 3D Gaussian Splatting with our method, showing a significant increase in PSNR.
3.2 Ablation studies
Blurriness loss and Event Integration Loss. To validate the effectiveness of training losses, we conduct the experiments on the "hotdog" scene from E2NeRF dataset. The results in the Tab.2 demonstrates that both proposed losses significantly improve the deblurring performance, with the event integration loss further slightly enhancing sharp rendering quality.
Table 2: Ablation study to validate the effectiveness of losses. "w" represents "with" and "wo" represents "without". "$L_{blur}$" denotes the blurriness loss function. "$L_{event}$" denotes the event integration loss.
[…]
PSNR 22.92 31.80 31.83
SSIM 0.886 0.9490 0.9488
With/Without MLP Latent Poses Estimation We conduct the experiments to further analyze the contribution of ADE module. Using initial latent poses estimated by COLMAP with EDI preprocessed images, we compare results with and without the ADE network. As shown in Table 3, the ADE module improves all evaluation metrics, indicating its effectiveness in estimating latent poses during exposure.
Table 3: Ablation study to analyze the contribution of ADE module. "w" represents "with" and "wo" represents "without".
Latent Pose Estimation wo MLP w MLP
PSNR 28.76 31.83
SSIM 0.9347 0.9488
Further ablations are provided in the supplementary materials.
4 Conclusion
In this paper, we introduce Event-Assisted 3D Deblur Gaussian Splatting (EaDeblur-GS), which integrates event data and RGB images to achieve sharp neural 3D representations. We propose an Adaptive Deviation Estimator network to estimate latent poses during exposure by computing Gaussian deviations. Our method, evaluated against state-of-the-art techniques and through extensive ablation studies, shows significant improvements over the original 3D Gaussian Splatting and performance comparable to Event-RGB deblur NeRF (E2NeRF).

### v2

@@
 To address these issues, we propose Event-assisted 3D Deblur Reconstruction with Gaussian Splatting (EaDeblur-GS), which integrates event camera data to enhance the robustness of 3DGS against motion blur.
-By employing an Adaptive Deviation Estimator (ADE) network to estimate Gaussian center deviations and using novel loss functions, EaDeblur-GS achieves sharp 3D reconstructions in real-time, demonstrating performance comparable to state-of-the-art methods.
+EaDeblur-GS utilizes a novel Adaptive Deviation Estimator (ADE) network and two novel loss functions to achieve real-time, sharp 3D reconstructions.
+Evaluations demonstrates that our method achieves advanced performance against original Gaussian Splatting and other Deblur Gaussian Splatting techniques.
 Keywords:
@@
 The advent of Neural Radiance Fields (NeRF) [4] has brought revolutionary advancements in photo-realistic novel view synthesis.
-Recently, The introduction of 3D Gaussian Splatting (3DGS) [1] has revolutionized 3D scene representation using Gaussian ellipsoids, achieving high-quality 2D rendering and faster speeds.
+Building on this progress, the recent introduction of 3D Gaussian Splatting (3DGS) [1] has further enhanced 3D scene representation using Gaussian ellipsoids, achieving high-quality 2D rendering and faster training and rendering speeds.
 However, in real-world applications, factors like camera shake and shutter speed often lead to image blurriness and inaccurate camera poses estimation, challenging the clear neural volumetric representation.
-Therefore, several methods have been developed to handle blurriness in NeRFs and 3DGS.
+Several methods address blurriness in NeRFs and 3DGS.
 NeRF’s deblurring technology has undergone relatively early development, with Deblur-NeRF [3] being the first framework to tackle this issue, which utilized an analysis-by-synthesis approach to recover sharp NeRFs from blurry inputs.
@@
 But the NeRF-based methods always consume extensive training time and rendering time.
-Some methods based on 3DGS has been developed because of its advantages in rendering and training speed.
+Hence, some methods based on 3DGS has been developed because of its advantages in rendering and training speed.
 For instance, Wenbo Chen et al.
@@
 [6] combined blurred images with event streams using innovative loss functions and the Event Double Integral (EDI) approach.
-For 3DGS, Yu proposed EvaGaussian[10], which enhanced 3DGS with assistance of event streams.
+For 3DGS, Yu proposed EvaGaussian[10], optimizing both scene reconstruction and camera trajectory estimation to achieve unparalleled detail and fidelity in real-time 3D scene synthesis.
 Despite achieving excellent performance in 3D deblurring reconstruction, the aforementioned methods still have limitations.
@@
 Our contributions are summarized as follows:
-  • We propose Event-assisted 3D Deblur Reconstruction with Gaussian Splatting (EaDeblur-GS), which incorporates blurry RGB images and event streams into Gaussian Splatting to recover sharp 3D representations.
+  • We propose Event-assisted 3D Deblur Gaussian Splatting (EaDeblur-GS), which incorporates blurry RGB images and event streams into Gaussian Splatting to recover sharp 3D representations.
   • We introduce a novel Adaptive Deviation Estimator (ADE) network to simulate the shaking motion during exposure accurately by estimating the deviations of Gaussians.
-  • We comprehensively evaluate the proposed method and compare it with several baselines, demonstrating that our EaDeblur-GS achieves performance comparable to state-of-the-art methods while enabling real-time sharp image rendering.
+  • We comprehensively evaluate the proposed method and compare it with several baselines, demonstrating that our EaDeblur-GS achieves advanced performance while enabling real-time sharp image rendering.
 2 Method
-As shown in Fig.1, our method starts with the input of blurry RGB images and the corresponding event streams.
-We first employ the Event Double Integral (EDI) technique [5] to generate a set of latent sharp images.
-These images are then processed by COLMAP to achieve enhanced initial reconstruction and precise camera pose estimation.
-From this enhanced reconstruction, we create a set of 3D Gaussians.
-Subsequently, the positions of these Gaussians and the estimated camera poses are fed into our proposed ADE network to determine the positional deviations of the Gaussians.
-These adjusted 3D Gaussians are then projected onto each viewpoint, including corresponding latent viewpoints, to produce sharp image renderings.
-Furthermore, we integrate a Blurriness Loss to simulate the generation of authentic blurry images, and an Event Integration Loss to guide the Gaussian model in accurately capturing the true shapes of objects.
-This allows the model to learn precise 3D volume representations and achieve superior 3D reconstructions.
-The overview of our method is illustrated in Fig.1.
+As illustrated in Fig.1, our method begins with the input of blurry RGB images and their corresponding event streams.
+We first use the Event Double Integral (EDI) technique to generate a set of latent sharp images.
+These images are then processed with COLMAP, enhancing the initial reconstruction and providing relative precise camera pose estimation.
+From this reconstruction, we create a set of 3D Gaussians.
+The positions of these Gaussians, along with the estimated camera poses, are input into our Adaptive Deviation Estimator (ADE) network to determine positional deviations, which are added to the original Gaussian centers.
+The adjusted 3D Gaussians are projected onto each viewpoint, including corresponding latent viewpoints, to render sharp images.
+We also introduce a Blurriness Loss to simulate realistic blurriness and an Event Integration Loss to improve object shape accuracy in the Gaussian model.
+This process enables the model to learn precise 3D volume representations and achieve superior 3D reconstructions.
+The overall method is depicted in Fig.1.
 Figure 1: The overview of EaDeblur-GS
@@
 $I_{k}=\frac{(b+1)I_{blur}e^{\Theta\sum_{i=1}^{k}B_{i}}}{1+e^{\Theta\sum_{i=1}^{1}B_{i}}+\cdots+e^{\Theta\sum_{i=1}^{b}B_{i}}}.$ (2)
-We then utilize COLMAP to estimate the poses of EDI-processed sharp images $\{\mathbf{P}_{k}\}_{k=0}^{b}=\mathrm{COLMAP}(\{I_{k}\}_{k=0}^{b})$, which also results in a more accurate initial sparse point cloud.
+We then utilize COLMAP to estimate the poses of EDI-processed sharp images ($(\{\mathbf{P}_{k}\}_{k=0}^{b}=\mathrm{COLMAP}(\{I_{k}\}_{k=0}^{b})$, which also results in a more accurate initial sparse point cloud.
 Inspired by [9] and [2], we represent latent poses as displacements of Gaussian ellipsoid centers $x_{j}$.
-To estimate the deviations of Gaussians in this context, we employ Adaptive Deviation Estimator network, which is consisted of a small Multi-Layer Perceptron (MLP) $\mathcal{F}\theta$ with three hidden linear layers and an embedding layer which transforms low-frequency positional information into high-frequency expressions.
-The ADE $\mathcal{F}\theta$ takes the EDI-predicted poses ${\mathbf{P}k}{k=0}^{b}$ and original positions of Gaussians $x_{j}$ to estimate deviations:
+To estimate the deviations of Gaussians in this context, we employ Adaptive Deviation Estimator network, which is consisted of a small Multi-Layer Perceptron (MLP) $\mathcal{F}_{\theta}$ with three hidden linear layers.
+Besides, we utilize an embedding layer which encodes low-frequency positional information into high-frequency expressions and a decode layer to output the estimated deviations.
+The ADE $\mathcal{F}_{\theta}$ takes the EDI-predicted poses ${\mathbf{P}}_{k=0}^{b}$ and original positions of Gaussians $x_{j}$ to estimate deviations:
 $\{(\delta x_{j}^{(i)}\}_{i=1}^{l}=\mathcal{F}_{\theta}\Big{(}\gamma(x_{j}),\gamma(P)\Big{)}$ (3)
 where $l$ means the number of estimated latent poses, and $\delta x_{j}^{(i)}$ denotes the $i$-th predicted position offset of $j$-th Gaussian.
-We then obtain extra $l$ sets of 3D Gaussians $\{\{\hat{x}_{j}^{(i)}\}_{i=1}^{l}\}_{j=1}^{N_{G}}$ by adjusting the positions of the original 3D Gaussians, where $N_{G}$ is the number of Gaussians, $\hat{x}_{j}^{(i)}$ represents the position offset scaled by $\lambda_{p}$: $\delta x_{j}^{(i)}$ by $\lambda_{p}.\hat{x}_{j}^{(i)}=x_{j}+\lambda_{p}\delta x_{j}^{(i)}$.
+We then obtain extra $l$ sets of 3D Gaussians $\{\{\hat{x}_{j}^{(i)}\}_{i=1}^{l}\}_{j=1}^{N_{G}}$ by adjusting the positions of the original 3D Gaussians, where $N_{G}$ is the number of Gaussians, $\hat{x}_{j}^{(i)}$ represents the position offset scaled by $\lambda_{p}:\hat{x}_{j}^{(i)}=x_{j}+\lambda_{p}\delta x_{j}^{(i)}$.
 This method computes $l$ different sets of 3D Gaussians for each viewpoint, which can be rasterized to $l$ sharp latent images $\{I_{i}\}_{i=0}^{l}$.
 Additionally, we rasterize an image with original Gaussians positions as the image rendered from original viewpoint.
-During forward rendering process, the MLP network and multiple rendering are not necessary, ensuring real-time inference speed comparable to original 3D Gaussian.
+During forward rendering process, the ADE network and multiple rendering are not necessary, ensuring real-time inference speed comparable to original 3D Gaussian.
 2.2 Loss Functions
@@
 $\hat{I}_{blur}=\frac{1}{b+1}\sum_{i=0}^{b}I_{i},\quad I_{i}=\text{Rasterize}(\{G(\hat{x}_{j}^{(i)})\}_{j=1}^{N_{G}})$ (4)
-where $I_{i}$ is a sharp image generated by the estimated corrections, $I_{blur}$ is the estimated blurry image.
+where $I_{i}$ is the sharp images generated by multiple rendering, $\hat{I}_{blur}$ is the estimated blurry image.
 The blurriness loss is computed as the difference between estimated blurry image and input blurry image $I_{blur}$, combined with a D-SSIM loss as follows:
@@
 Leveraging the high time-resolution event stream, we adopt an Event Integration Loss to guide the network in learning a fine-grained sharp 3D reconstructions.
-We integrate the events polarities between two input frames as follows:
+We integrate the events polarities between two input blurry frames as follows:
 […]
@@
 3.1 Comparisons and Results
-To evaluate the effectiveness of using event data to aid in deblurring RGB images, we compare our method with the original Gaussian Splatting (GS), which utilizes blurry images as input and supervision.
+To evaluate the effectiveness of the proposed method, we compare our method with the original Gaussian Splatting (GS), which utilizes blurry images as input and supervision.
 Since COLMAP fails with only blurry images, we estimate the initial point cloud using EDI-deblurred images.
 We also compare our method with Deblurring 3D Gaussian Splatting(Deblurring-GS) to highlight the advantages of incorporating event data.
-Additionally, we benchmark our approach against Event Enhanced Neural Radiance Fields from Blurry Images(E2NeRF), the state-of-the-art method that leverages event data for deblurring RGB images.
 Our evaluation metrics include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Frames Per Second (FPS).
-All methods are tested on the synthetic data from E2NeRF, which includes perfectly sharp images from the same viewpoint as blurry input images with corresponding event streams.
+All methods are tested on the synthetic data from E2NeRF.
 The implementation details are provided in the supplementary materials.
 The quantitative comparison results are presented in Table 1 and the qualitative comparison is also provided in supplementary materials.
-Table 1: Quantitative analysis. The results in tha table are the averages of six synthetic scenes from E2NeRF. We use bold to mark the best result.
-Image Deblur GS Deblurring-GS E2NeRF Ours
-PSNR 22.15 22.70 29.77 30.08
-SSIM 0.878 0.8427 0.960 0.9366
-FPS 160 160 2 160
-Our results demonstrate that our method achieves performance comparable to the current state-of-the-art model in terms of PSNR and SSIM, while maintaining real-time rendering capabilities with noticeable FPS.
-The advantage of using event data for deblurring RGB images is further evidenced by comparing Deblurring 3D Gaussian Splatting with our method, showing a significant increase in PSNR.
+Table 1: Quantitative analysis. The results in the table are the averages of six synthetic scenes from E2NeRF. We use bold to mark the best result.
+Image Deblur GS Deblurring-GS Ours
+PSNR 22.15 22.70 29.69
+SSIM 0.878 0.8427 0.9338
+FPS 190 190 190
+The results demonstrate that our method achieves advanced performance compared with other Gaussian Splatting techniques in terms of PSNR and SSIM, while maintaining real-time rendering capabilities with noticeable FPS.
+The advantage of using event data for deblurring RGB images is further evidenced by comparing Deblurring 3D Gaussian Splatting with our method, showing a significant increase in PSNR and SSIM.
 3.2 Ablation studies
@@
 SSIM 0.886 0.9490 0.9488
-With/Without MLP Latent Poses Estimation We conduct the experiments to further analyze the contribution of ADE module.
+With/Without ADE network.
+We conduct the experiments to further analyze the contribution of ADE module.
 Using initial latent poses estimated by COLMAP with EDI preprocessed images, we compare results with and without the ADE network.
@@
 4 Conclusion
-In this paper, we introduce Event-Assisted 3D Deblur Gaussian Splatting (EaDeblur-GS), which integrates event data and RGB images to achieve sharp neural 3D representations.
+In this paper, we introduce EaDeblur-GS, which integrates event data and RGB images to achieve sharp neural 3D representations.
 We propose an Adaptive Deviation Estimator network to estimate latent poses during exposure by computing Gaussian deviations.
-Our method, evaluated against state-of-the-art techniques and through extensive ablation studies, shows significant improvements over the original 3D Gaussian Splatting and performance comparable to Event-RGB deblur NeRF (E2NeRF).
+Our method, evaluated against other techniques and through extensive ablation studies, shows significant improvements over the original 3D Gaussian Splatting and Deblurring 3D Gaussian Splatting.

### v3

@@
 From this reconstruction, we create a set of 3D Gaussians.
-The positions of these Gaussians, along with the estimated camera poses, are input into our Adaptive Deviation Estimator (ADE) network to determine positional deviations, which are added to the original Gaussian centers.
+These Gaussian positions, combined with the estimated camera poses, are then fed into our Adaptive Deviation Estimator (ADE) network to adjust the Gaussian centers by calculating positional deviations.
 The adjusted 3D Gaussians are projected onto each viewpoint, including corresponding latent viewpoints, to render sharp images.
@@
 The overall method is depicted in Fig.1.
-Figure 1: The overview of EaDeblur-GS
+Figure 1: The Overview of our approach.
+Our approach integrates blurred RGB images with event data to enhance image clarity.
+The EDI technique is applied to produce latent sharp images, which are then refined through COLMAP for accurate SFM reconstruction and 3D Gaussian modeling.
+The ADE module estimates positional deviations based on initial Gaussian positions and camera extrinsics, simulating camera motion.
+These deviated Gaussians are rendered into multiple views and compute blurriness and event integration losses, facilitating the learning of detailed 3D representations for enhanced reconstruction quality.
 In the following sections, we detail how the ADE network estimates deviations, followed by an in-depth introduction to blurriness loss and event integration loss.
@@
 The implementation details are provided in the supplementary materials.
-The quantitative comparison results are presented in Table 1 and the qualitative comparison is also provided in supplementary materials.
+The quantitative comparison results are presented in Table 1.
 Table 1: Quantitative analysis. The results in the table are the averages of six synthetic scenes from E2NeRF. We use bold to mark the best result.
@@
 The advantage of using event data for deblurring RGB images is further evidenced by comparing Deblurring 3D Gaussian Splatting with our method, showing a significant increase in PSNR and SSIM.
+Figure 2: Qualitative results on E2NeRF synthetic dataset.
+We also assess the qualitative performance of our method compared to other approaches.
+As illustrated in Fig.2, the original 3D Gaussian Splatting struggles to produce clear 3D representations in certain scenes, even with EDI-preprocessed relatively sharp input images.
+While the Deblurring 3D Gaussian Splatting method can reconstruct relatively sharp objects, it still fails to capture fine details in some cases, such as with the “materials” object.
+In contrast, our method demonstrates the capability to reconstruct fine-grained, clear objects with greater accuracy.
 3.2 Ablation studies
 Blurriness loss and Event Integration Loss.
-To validate the effectiveness of training losses, we conduct the experiments on the "hotdog" scene from E2NeRF dataset.
+To validate the effectiveness of training losses, we conduct the experiments on the “hotdog" scene from E2NeRF dataset.
 The results in the Tab.2 demonstrates that both proposed losses significantly improve the deblurring performance, with the event integration loss further slightly enhancing sharp rendering quality.
@@
 Our method, evaluated against other techniques and through extensive ablation studies, shows significant improvements over the original 3D Gaussian Splatting and Deblurring 3D Gaussian Splatting.
+Acknowledgments
+This work was supported by the National Science and Technology Major Project under Grant 2020AAA0107300, and the Key Research and Development Plan Project of Xinjiang Uygur Autonomous Region under Grant 2023B01006-1.